# A Plug-and-Play Method for Guided Multi-contrast MRI Reconstruction based on Content/Style Modeling

### v2

@@ II-C Combining Image Translation with Reconstruction
Acknowledging the limitation of MR cross-contrast prediction, some prior work has attempted combining it with multi-contrast reconstruction. A naive form of joint image synthesis and reconstruction, e.g. PROSIT [39] and rsGAN [40], involves generating a synthetic image from the reference scan via deterministic image translation and using it in a classical L2-regularized least-squares reconstruction. More recent work by Xuan et al. [41] proposed a joint image translation and reconstruction method which additionally accounts for misalignment between the reference and reconstructing images, with a follow-up work leveraging optimal transport theory [42]. Levac et al. [16] formulate guided reconstruction as a highly general Bayesian maximum a posteriori estimation problem and solve it iteratively via Langevin update steps, using an image-domain diffusion model as the score function of the prior distribution. While these are promising directions, we propose an alternative approach that decomposes the multi-contrast problem in a more intuitive way – first, into two sub-problems, namely image modeling and image reconstruction; and second, within the image-domain model, the multi-contrast information is decomposed into content and style. This two-level decomposition results in a highly modular reconstruction algorithm with a built-in explanatory framework where (a) the guidance mechanism is a simple content-replacement operation, (b) the discrepancy between the supplied reference content and the true content of the target image represents a meaningful error term which can be minimized, and (c) the optimal content-encoding capacity of the model for a given two-contrast image dataset indicates the amount of shared structure available to be learned in this data and used in the reconstruction task.
III Methods
III-A Reconstruction Problem
III-A1 Undersampled MRI reconstruction
@@
III-A2 Plug-and-play denoiser
Plug-and-play methods replace the analytical operation of (2) with off-the-shelf denoisers. A CNN-based denoiser is of special interest as it incorporates a learning-based component into iterative reconstruction. Given a CNN model $f_{\theta}$ trained to remove i.i.d. Gaussian noise from an image, PnP-CNN [23] modifies Eq. (2) to
$r^{k}\leftarrow f_{\theta}(x^{k-1}).$ (4)
III-A3 Plug-and-play content consistency operator
@@
$r_{2}^{k}\leftarrow g_{M}(x_{2}^{k-1};c),$ (5)
where a content/style model $M$ decomposes $x_{2}^{k-1}$ into content and style, followed by a replacement of this corrupted content with high-quality content $c$ derived from $x^{\mathrm{ref}}_{1}$ and composing the improved image $r_{2}^{k}$ from it. Before formally defining this operator and developing the reconstruction algorithm, we discuss the steps required to learn the content/style model $M$.
III-B Content/Style Modeling
Given an image dataset of two MR contrasts we make the general assumption that there exists an underlying contrast-independent structure which, influenced by arbitrary contrast-specific factors, is rendered as the contrast images.
III-B1 Unpaired pre-training of MUNIT
We formulate our content/style model based on the MUNIT framework [22]. We define an image domain as the set $\mathcal{X}_{i}$ of images of a certain contrast comprising the dataset, where $i\in\{1,2\}$. “Content” $c\in\mathcal{C}$ is defined here as the underlying contrast-independent structure and is represented as a set of feature maps, whereas “style” $s_{i}\in\mathcal{S}_{i}$ corresponds to the various modes of variability in one domain which cannot be explained by the other, e.g. global effects of acquisition settings, contrast-specific tissue features, etc., and is represented as a low-dimensional vector. MUNIT posits the existence of functions $G_{i}^{*}:\mathcal{C}\times\mathcal{S}_{i}\rightarrow\mathcal{X}_{i}$ and their inverses $E_{i}^{*}=(G_{i}^{*})^{-1}$, and learns them jointly via unpaired training, given samples from marginal distributions $p(x_{i})$. In practice, the encoder $E_{i}$ is split into content encoder $E_{i}^{c}$ and style encoder $E_{i}^{s}$. Thus, the content/style model is specified as $M=\{E_{1}^{c},E_{2}^{c},E_{1}^{s},E_{2}^{s},G_{1},G_{2}\}$.
The MUNIT loss function is comprised of 4 terms:
@@ III-B1 Unpaired pre-training of MUNIT
$\mathcal{L}_{\mathrm{image}}^{\mathrm{self}}$ is the image recovery loss, which promotes preservation of the image information in the latent space. $\mathcal{L}_{\mathrm{content}}^{\mathrm{self}}$ and $\mathcal{L}_{\mathrm{style}}^{\mathrm{self}}$ are the content and style recovery losses, respectively, which specialize parts of the latent code into content and style components. $\mathcal{L}_{\mathrm{GAN}}$ is the adversarial loss, which enables unpaired training by enforcing distribution-level similarity between synthetic and real images via discriminators $D_{1}$ and $D_{2}$. $\alpha_{1}$, $\alpha_{2}$, and $\alpha_{3}$ are hyperparameters.
III-B2 Paired fine-tuning
While the pre-training stage learns useful content/style representations, the model can be adapted to our task by improving its content preservation. To this end, we propose a paired fine-tuning (PFT) stage, leveraging a modest amount of paired data. Given samples from the joint distribution $p(x_{1},x_{2})$, our fine-tuning objective is given as
@@ III-B2 Paired fine-tuning
$\begin{split}\mathcal{L}_{\mathrm{image}}^{\mathrm{cross}}=\>&\mathbb{E}_{\{x_{1},x_{2}\}\sim p(x_{1},x_{2})}[\\ &||x_{2}-G_{2}(E_{1}^{c}(x_{1}),E_{2}^{s}(x_{2}))||_{1}+\\ &||x_{1}-G_{1}(E_{2}^{c}(x_{2}),E_{1}^{s}(x_{1}))||_{1}],\end{split}$ (12)
@@
$\mathcal{L}_{\mathrm{image}}^{\mathrm{cross}}$ is a pixel-wise image translation loss, which provides image-level supervision, and $\mathcal{L}_{\mathrm{content}}^{\mathrm{cross}}$ is a paired content loss, which penalizes discrepancy between the contents. $\beta_{1}$, $\beta_{2}$, and $\beta_{3}$ are an additional set of hyperparameters.
III-B3 Network architecture and content capacity
Following the original paper [22], our content encoders $E_{i}^{c}$ consist of an input convolutional layer potentially followed by strided downsampling convolutions, and finally a series of residual blocks. Style encoders $E_{i}^{s}$ consist of input and downsampling convolutions followed by adaptive average pooling and a fully-connected layer that outputs the latent vector. Decoders $G_{i}$ follow a similar structure as the content encoders except in reverse. Style is introduced into the decoder via AdaIN operations which modulate the activation maps derived from the content. We observe that in this architecture, the ratio between the content and image resolutions reflects the level of local structure one expects to be shared between the two domains. This (relative) content resolution is thus an inductive bias built into the model, which we refer to as the model’s content capacity. This concept is closely related to the general concept of the locality bias of image-to-image models [43]. A model with high content capacity has a large content resolution, allowing it to learn a rich content representation that strongly influences the output’s structure, which simultaneously restricts the style spaces to learn low-level global features.
Figure 2: A trained content/style model $M$ and the content consistency operator $g_{M}(\cdot;c)$ based on it. Images shown here are from the NYU brain DICOM dataset. Practical details are provided in Section IV. (a) A T1W/T2W image pair and the corresponding content maps. While these two contents generally agree, there is a notable discrepancy. (b) Synthetic T2W images generated from 16 style codes grid-sampled from $\mathcal{S}_{2}$, showing a smooth variation in image contrast and a roughly constant anatomical structure. (c) $g_{M}(\cdot;c)$ is applied on 3 cases where $x_{2}^{*}$ was corrupted with $R$=4 undersampling with different center fractions. Sampling more low-frequency lines leads to more accurate style estimate $\hat{s}_{2}$ and thus a better image $x_{2}^{\mathrm{cc}}$. Contours in $\mathcal{S}_{2}$ indicate the MAE of the synthetic images from that region. The red circle indicate a style estimation NMSE of 0.1.
III-C Iterative Reconstruction using Content Consistency
@@
Note that this is a radical operation as it discards all structure contained in $x^{\mathrm{us}}_{2}$, retaining only a compact style code $\hat{s}_{2}=E_{2}^{s}(x^{\mathrm{us}}_{2})$. Let $x_{2}^{*}$ be the ground truth reconstruction with content $c^{*}$ and style $s_{2}^{*}$. The success of our content consistency operation depends on two conditions – (a) $\hat{c}$ is close to $c^{*}$ and (b) $\hat{s}_{2}$ is close to $s_{2}^{*}$. The first condition is roughly satisfied, as seen in Fig. 2a, because the model is explicitly trained to minimize content discrepancy, but more on this later. On the other hand, it is not obvious that the second condition should hold too and hence deserves a closer look. Assuming a high degree of shared local structure between the two domains, the optimal model has high content capacity (see Section III-B3). Thus, style would represent mostly low-level global image features, e.g. contrast variations, as indeed observe in Fig. 2b. It is a well known fact that image contrast is contained prominently in the center of the k-space. Hence, the estimate $\hat{s}_{2}$ can be made arbitrarily close to $s_{2}^{*}$ by sufficiently sampling the k-space center, as shown empirically in Fig. 2c.
@@
Algorithm 1 PnP-MUNIT
@@
7:     $r_{2}^{k}\leftarrow g_{M}(x_{2}^{k-1};c^{k-1})$triangleright Content consistency
8:     $x_{2}^{k}\leftarrow r_{2}^{k}-\eta A^{H}(Ar_{2}^{k}-y)$triangleright Data consistency
9:     $c^{k}\leftarrow c^{k-1}-\gamma\nabla_{c}||AG_{2}(c^{k-1},E_{2}^{s}(x_{2}^{k}))-y||_{2}^{2}$triangleright CR
@@
The core assumption of our idealized content/style model is that contents $\hat{c}$ and $c^{*}$ are identical. However, in reality, this assumption will not hold, and a discrepancy between the two contents is to be expected. There are two possible sources of this discrepancy – (a) model-related, e.g. fundamental limits such as irreducible error [^1] If perfectly zero error was possible, one could perfectly predict the (distribution of) target image from the reference without any measurement. and practical issues like sub-optimal training, and (b) reference image-related, e.g. presence of artifacts independent of the target reconstruction. This content discrepancy, as observed empirically, is shown in Fig. 2a. During reconstruction, the error in $\hat{c}$ would limit the efficacy of the operator $g_{M}(\cdot;\hat{c})$, affecting reconstruction quality. While model-related discrepancy is partly tackled by PFT (Section III-B2), we now propose a content refinement (CR) process to correct for the remaining discrepancy in the reconstruction stage. Since we have no direct access to the true content $c^{*}$, but only to undersampled k-space measurements $y$, we aim at solving the following minimization problem in CR
@@
initialized as $c^{0}\leftarrow E^{c}_{1}(x^{\mathrm{ref}}_{1})$ and updated every $k^{\text{th}}$ iteration with step size $\gamma$ following the data consistency and content consistency updates. Hence, by aligning the content $c$ with k-space data $y$ and correcting the discrepancy, the CR module aligns content consistency updates with data consistency updates. With this additional component in place, we obtain our PnP-MUNIT reconstruction algorithm (Algorithm 1). Note that unlike content, the style of the image estimate need not be explicitly aligned with the k-space. The data consistency update applied on the image implicitly “corrects” its style and as the image converges, its style would converge as well.
IV Experimental Setup
IV-A Experiment Design
@@
Without loss of generality, we considered the case of reconstructing T2W scans using T1W references and focused on head applications. We conducted 4 sets of experiments to comprehensively evaluate PnP-MUNIT – (a) simulations, (b) benchmark on NYU DICOM data, (c) benchmark on clinical multi-coil raw data, and (d) radiological evaluation.
TABLE I: Overview of the sequences in the LUMC datasets.

### v3

@@ 2.3 Combining image translation with reconstruction
Acknowledging the limitation of MR cross-contrast prediction, some prior work has attempted combining it with multi-contrast reconstruction. A naive form of joint image synthesis and reconstruction, e.g. PROSIT (Mattern et al., 2020) and rsGAN (Dar et al., 2020), involves generating a synthetic image from the reference scan via deterministic image translation and using it in a classical L2-regularized least-squares reconstruction. More recent work by Xuan et al. (2022) proposed a joint image translation and reconstruction method which additionally accounts for misalignment between the reference and reconstructing images, with a follow-up work leveraging optimal transport theory (Wang et al., 2024). Levac et al. (2023) formulate guided reconstruction as a highly general Bayesian maximum a posteriori estimation problem and solve it iteratively via Langevin update steps, using an image-domain diffusion model as the score function of the prior distribution. While these are promising directions, we propose an alternative approach that decomposes the multi-contrast problem in a more intuitive way – first, into two sub-problems, namely image modeling and image reconstruction; and second, within the image-domain model, the multi-contrast information is decomposed into content and style. This two-level decomposition results in a highly modular reconstruction algorithm with a built-in explanatory framework where (a) the guidance mechanism is a simple content-replacement operation, (b) the discrepancy between the supplied reference content and the true content of the target image represents a meaningful error term which can be minimized, and (c) the optimal content-encoding capacity of the model for a given two-contrast image dataset indicates the amount of shared structure that is available to be learned in this data and utilized in the reconstruction task.
3 Methods
3.1 Reconstruction problem
3.1.1 Undersampled MRI reconstruction
@@
3.1.2 Plug-and-play denoiser
Plug-and-play methods replace the analytical operation of (2) with off-the-shelf denoisers. A CNN-based denoiser is of special interest as it incorporates a learning-based component into iterative reconstruction. Given a CNN model $f_{\theta}$ trained to remove i.i.d. Gaussian noise from an image, PnP-CNN (Ahmad et al., 2020) modifies Eq. (2) to
$z^{k}\leftarrow f_{\theta}(x^{k-1}).$ (4)
3.1.3 Plug-and-play content consistency operator
@@
$z_{2}^{k}\leftarrow g_{M}(x_{2}^{k-1};c),$ (5)
where a content/style model $M$ decomposes $x_{2}^{k-1}$ into content and style, followed by a replacement of this corrupted content with high-quality content $c$ derived from $x^{\mathrm{ref}}_{1}$ and composing the improved image $z_{2}^{k}$ from it. Before formally defining this operator and developing the reconstruction algorithm, we discuss in the following section the design and learning of the content/style model $M$.
3.2 Content/Style modeling
Given an image dataset of two MR contrasts we make the general assumption that there exists an underlying contrast-independent structure which, influenced by arbitrary contrast-specific factors, is rendered as the contrast images. We formulate our content/style model based on the MUNIT framework (Huang et al., 2018).
3.2.1 Unpaired MUNIT pre-training
We define an image domain as the set $\mathcal{X}_{i}$ of images of a certain contrast comprising the dataset, where $i\in\{1,2\}$. ”Content” $c\in\mathcal{C}$ is defined as the underlying contrast-independent structure and is represented as a set of feature maps, whereas ”style” $s_{i}\in\mathcal{S}_{i}$ corresponds to the various modes of variability in one domain which cannot be explained by the other, e.g. global effects of acquisition settings, contrast-specific tissue features, etc., and is represented as a low-dimensional vector. MUNIT posits the existence of functions $G_{i}^{*}:\mathcal{C}\times\mathcal{S}_{i}\rightarrow\mathcal{X}_{i}$ and their inverses $E_{i}^{*}=(G_{i}^{*})^{-1}$, and learns them jointly via unpaired training, given samples from marginal distributions $p(x_{i})$. In practice, the encoder $E_{i}$ is split into content encoder $E_{i}^{c}$ and style encoder $E_{i}^{s}$. Thus, the content/style model is specified as $M=\{E_{1}^{c},E_{2}^{c},E_{1}^{s},E_{2}^{s},G_{1},G_{2}\}$.
The MUNIT loss function is comprised of four terms:
@@ 3.2.1 Unpaired MUNIT pre-training
$\mathcal{L}_{\mathrm{image}}^{\mathrm{self}}$ is the image recovery loss, which promotes preservation of the image information in the latent space. $\mathcal{L}_{\mathrm{content}}^{\mathrm{self}}$ and $\mathcal{L}_{\mathrm{style}}^{\mathrm{self}}$ are the content and style recovery losses, respectively, which specialize parts of the latent code into content and style components. $\mathcal{L}_{\mathrm{GAN}}$ is the adversarial loss, which enables unpaired training by enforcing distribution-level similarity between synthetic and real images via discriminators $D_{1}$ and $D_{2}$. $\alpha_{1}$, $\alpha_{2}$, and $\alpha_{3}$ are hyperparameters. For our purpose, we consider the MUNIT training process as the pre-training stage of our content/style model.
3.2.2 Paired fine-tuning
Figure 2: Our proposed paired fine-tuning (PFT) stage for improving the alignment of the content representations of the two domains given a small dataset of paired images.
While the pre-training stage learns useful content/style representations, we adapt the model to our task by improving its content preservation. To this end, we propose a paired fine-tuning (PFT) stage which more strongly aligns the content representations of the two domains given a modest amount of paired data. Given samples from the joint distribution $p(x_{1},x_{2})$, our fine-tuning objective is expressed as
@@ 3.2.2 Paired fine-tuning
$\begin{split}\mathcal{L}_{\mathrm{image}}^{\mathrm{cross}}=\mathbb{E}_{\{x_{1},x_{2}\}\sim p(x_{1},x_{2})}[\>&||x_{2}-G_{2}(E_{1}^{c}(x_{1}),E_{2}^{s}(x_{2}))||_{1}+\\ &||x_{1}-G_{1}(E_{2}^{c}(x_{2}),E_{1}^{s}(x_{1}))||_{1}],\end{split}$ (12)
@@
$\mathcal{L}_{\mathrm{image}}^{\mathrm{cross}}$ is a pixel-wise image translation loss, which provides image-level supervision, and $\mathcal{L}_{\mathrm{content}}^{\mathrm{cross}}$ is a paired content loss, which penalizes discrepancy between the contents. $\beta_{1}$, $\beta_{2}$, and $\beta_{3}$ are an additional set of hyperparameters. Note that our PFT preserves the bi-directionality of the content/style model, a property which will later be essential for generalization over different reconstruction tasks.
3.2.3 Network architecture and content capacity
Following the original MUNIT paper (Huang et al., 2018), our content encoders $E_{i}^{c}$ consist of an input convolutional layer potentially followed by strided downsampling convolutions, and finally a series of residual blocks. Style encoders $E_{i}^{s}$ consist of input and downsampling convolutions followed by adaptive average pooling and a fully-connected layer that outputs the latent vector. Decoders $G_{i}$ follow a similar structure as the content encoders except in reverse. Style is introduced into the decoder via AdaIN operations which modulate the activation maps derived from the content. We observe that in this architecture, the ratio between the content resolution and the image resolution reflects the level of local structure one expects to be shared between the two domains. This (relative) content resolution is thus an inductive bias built into the model, which we refer to as the model’s content capacity and define as
$J_{M}:=\frac{H_{c}W_{c}}{H_{x}W_{x}},$ (14)
where $H_{x}$ and $W_{x}$ are the height and width of the input image and $H_{c}$ ($\leq H_{x}$) and $W_{c}$ ($\leq W_{x}$) are the those of the corresponding content maps. This ratio is controlled via the number of up/downsampling blocks in the networks. Our notion of content capacity is closely related to the more general concept of the locality bias of image-to-image models (Richardson and Weiss, 2021). A content/style model with a high content capacity has a large content resolution, allowing it to learn a rich, highly expressive content representation that strongly influences the output’s structure. Consequently. the model’s style spaces are restricted to learn mostly low-level global features.
Figure 3: Demonstration of the content consistency operator $g_{M}(\cdot;c)$ based on a content/style model $M$. The image pair shown here are NYU brain DICOMs, practical details regarding which are provided in Section 4. (a) A T1W/T2W image pair and the corresponding content maps. While these two content maps generally agree, there is a notable discrepancy. (b) Synthetic T2W images generated from 16 style codes grid-sampled from $\mathcal{S}_{2}$, showing a smooth variation in image contrast and a roughly constant anatomical structure. (c) $g_{M}(\cdot;c)$ is applied on 3 cases where $x_{2}^{*}$ was corrupted with $R$=4 undersampling with different center fractions. Sampling more low-frequency lines leads to more accurate style estimate $\hat{s}_{2}$ and thus a better image $x_{2}^{\mathrm{cc}}$. Contours in $\mathcal{S}_{2}$ indicate the MAE of the synthetic images from that region. The red circle indicates a style estimation NMSE of 0.1.
3.3 Iterative reconstruction based on content consistency
@@
Note that this is a radical operation as it discards all structure contained in $x^{\mathrm{us}}_{2}$, retaining only a compact style code $\hat{s}_{2}=E_{2}^{s}(x^{\mathrm{us}}_{2})$. Let $x_{2}^{*}$ be the ground truth reconstruction with content $c^{*}$ and style $s_{2}^{*}$. The success of our content consistency operation depends on three conditions – (a) $\hat{c}$ and $c^{*}$ are rich, high-resolution representations of the underlying anatomy, (b) $\hat{c}$ is close to $c^{*}$, and (c) $\hat{s}_{2}$ is close to $s_{2}^{*}$. The first condition is satisfied provided that there is sufficient amount of shared structure between the reference and target domains and that the model has high content capacity $J_{M}$ for capturing a highly expressive content. The second condition is roughly satisfied, as seen in Fig. 3a, because the model is explicitly trained to minimize content error, but more on this later. On the other hand, it is not obvious that the third condition should hold too and it hence deserves a closer look. If the first condition is satisfied, the optimal model will have a high content capacity and consequently, style would represent mostly low-level global image features, e.g. contrast variations (see Section 3.2.3). This is observed in practice as shown in Fig. 3b. It is a well known fact that image contrast is contained prominently in the center of the k-space. Hence, the estimate $\hat{s}_{2}$ can be made arbitrarily close to $s_{2}^{*}$ by sufficiently sampling the k-space center, as demonstrated empirically in Fig. 3c.
@@
Algorithm 1 PnP-CoSMo iterative guided reconstruction
@@
7:     $z_{2}^{k}\leftarrow g_{M}(x_{2}^{k-1};c^{k-1})$triangleright Content consistency update
8:     $x_{2}^{k}\leftarrow z_{2}^{k}-\eta A^{H}(Az_{2}^{k}-y)$triangleright Data consistency update
9:     $c^{k}\leftarrow c^{k-1}-\gamma\nabla_{c}||AG_{2}(c^{k-1},E_{2}^{s}(x_{2}^{k}))-y||_{2}^{2}$triangleright Content refinement (CR) update
@@
The core assumption of an idealized content/style model is that contents $\hat{c}$ and $c^{*}$ are identical. However, in reality, this assumption will not hold, and a discrepancy between the two contents is to be expected. There are two possible sources of this discrepancy – (a) model-related, e.g. fundamental limits such as irreducible error [^1] If perfectly zero error was possible, one could perfectly predict the (distribution of) target image from the reference without any measurement. and practical issues like sub-optimal network design and training, and (b) reference image-related, e.g. presence of artifacts independent of the target reconstruction. We refer to this gap between $\hat{c}$ and $c^{*}$ as content discrepancy. Fig.3a shows content discrepancy as observed empirically. During reconstruction, the error in $\hat{c}$ would limit the efficacy of the operator $g_{M}(\cdot;\hat{c})$, affecting the reconstruction quality. While model-related discrepancy is partly tackled by PFT (Section 3.2.2), we now propose a content refinement (CR) process to correct for the remaining discrepancy in the reconstruction stage. Since we have no direct access to the true content $c^{*}$, but only to undersampled k-space measurements $y$, we aim at solving the following minimization problem in CR
@@
initialized as $c^{0}\leftarrow E^{c}_{1}(x^{\mathrm{ref}}_{1})$ and updated every $k^{\text{th}}$ iteration with step size $\gamma$ following the data consistency and content consistency updates. Hence, by aligning the content $c$ with k-space data $y$ and correcting the discrepancy, the CR module aligns content consistency updates with data consistency updates. With this additional component in place, we obtain our PnP-CoSMo reconstruction algorithm (Algorithm 1). Note that unlike content, the style of the image estimate need not be explicitly aligned with the k-space. Since the style information is assumed to be derived largely from the k-space data, the data consistency update applied on the image implicitly ”corrects” its style. Thus, as the image converges, we hypothesize that style would converge as well.
4 Experimental setup
4.1 Experiment design
@@
We conducted 4 sets of experiments to comprehensively evaluate PnP-CoSMo – (a) simulations, (b) benchmark against end-to-end methods on NYU DICOM data, (c) benchmark and ablation on clinical multi-coil raw data, and (d) radiological evaluation. Without loss of generality, we considered the case of reconstructing T2W scans using T1W references and focused on head applications, except in the NYU benchmark where the reverse case was considered as well.
4.1.1 Simulation
The goal of our simulation experiments was to empirically test several properties of our algorithm under highly controlled settings. To this end, we used simulated T1W and T2W images based on BrainWeb phantoms (Collins et al., 1998). BrainWeb provides 20 anatomical models of normal brain, each comprised of fuzzy segmentation maps of 12 tissue types. The 20 volumes were first split into the model training (18), model validation (1), and reconstruction testing (1) subsets. T1W/T2W spin-echo scans were simulated using TE/TR values randomly sampled from realistic ranges. For testing the reconstruction, 2D single-coil T2W k-space data was simulated via Fourier transform and 1D Cartesian random sampling at various accelerations. We analyzed the effect of the content/style model’s disentanglement level and its content capacity on the reconstruction quality, the convergence of the PnP-CoSMo algorithm, and the effectiveness, robustness, and tuning of the CR module.
4.1.2 Benchmark on the NYU DICOM dataset
In the NYU benchmark, we compared PnP-CoSMo against end-to-end reconstruction methods. The question we sought to answer was whether or not PnP-CoSMo, which requires only image-domain training data, can outperform methods that additionally require k-space training data. Furthermore, end-to-end methods are task-specific where the reference and target contrast are fixed – e.g. T1W and T2W, respectively. On the other hand, the content/style model used in PnP-CoSMo does not possess this kind of uni-directionality. Hence, we sought to assess the generalizability of PnP-CoSMo across two different reconstruction tasks, namely T1W-guided T2W reconstruction and T2W-guided T1W reconstruction. The NYU brain dataset (Knoll et al., 2020; Zbontar et al., 1811) includes raw data and DICOM scans of 4 contrasts – T1W with and without gadolinium agent, T2W, and FLAIR. For multi-contrast experiments, however, the NYU raw dataset does not provide subject labels for pairing the contrasts. Therefore, instead of the NYU raw data, we used the NYU DICOMs in this study, following Xuan et al. (2022). A paired DICOM subset of 327 subjects was obtained based on T2W and non-gadolinium T1W scans. All T1W scans were rigid-registered with the corresponding T2W scans. Single-coil k-space data was simulated via Fourier transform with 1D Cartesian random sampling at $R\in\{2,3,4,5\}$ and added Gaussian noise of $\sigma=0.01\mathrm{max}(x_{2}^{*})$. The dataset was split into 4 subsets – model-train (200), model-val (27), recon-val (50), and recon-test (50). Images from the former two splits were used to pre-train, fine-tune, and validate the content/style model. End-to-end models were trained and validated on the aligned T1W, the simulated T2W k-space, and the T2W ground-truth from these two splits. The recon-val split was used to tune the PnP-CoSMo algorithm, and recon-test was the held-out test set. We benchmarked PnP-CoSMo against one well-known single-contrast network – MoDL (Aggarwal et al., 2018) – and two recent multi-contrast networks – MTrans (Feng et al., 2022) and MC-VarNet (Lei et al., 2023). These networks were evaluated on the aforementioned two reconstruction tasks, assessing both their in-distribution and out-of-distribution performance – e.g. for the T1W-guided T2W reconstruction task, in-distribution means that the end-to-end model was trained on this task, whereas out-of-distribution means the model was trained of the T2W-guided T1W reconstruction task.
Table 1: Overview of the sequences in the LUMC datasets.